OpenStudy (anonymous):

Is the slope of the line through the points (1, -3) and (3, -5) = 1?

OpenStudy (anonymous):

(-5-(-3))/(3-1) = -2/2=-1

OpenStudy (anonymous):

use the slope formula when given 2 points

OpenStudy (lalaly):

\[\huge{slope}=\large{\frac{-5--3}{3-1}}=\frac{-2}{2}=-1\]